Circular No. 8161

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003gi AND 2003gj Additional apparent supernovae found on unfiltered CCD images have

been reported by L. A. G. Monard (SN 2003gi; cf. IAUC 8159) and by M. Moore and W. Li (SN 2003gj; LOTOSS/KAIT; cf. IAUC 8158):

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003gi May 17.108 0^{h}38^{m}32.^{s}67 −24^{o}20′33′′.5 15.6 1′′ E, 9′′ S
2003gj June 30.4 21 07 19.18 −25 29 23.6 17.7 14′′.5 W, 8′′.9 S

Additional approximate magnitudes from the respective observers: SN 2003gi in IC 1561, Apr. 29.134 UT, [18.0; July 1.103, 16.2. SN 2003gj in NGC 7017, June 3.4, [18.5; July 1.4, 17.6. Nothing is present at the location of SN 2003gi on the Digitized Sky Survey (limiting red mag 19). The offset for SN 2003gj is with respect to the western component of the interacting galaxy pair NGC 7017.

SUPERNOVA 2003gh IN NGC 2466 Corrigendum. On IAUC 8159, line 4, for 6′′ west read 6′′ east

SUPERNOVA 2002ic K. S. Kawabata, Y. Ohyama, and the FOCAS team, National Astro-

nomical Observatory of Japan; K. Nomoto, University of Tokyo; P. Mazzali, Trieste Observatory; and L. Wang, Lawrence Berkeley National Laboratory, write: “We obtained optical spectropolarimetry (range 380–1020 nm) of SN 2002ic (cf. IAUC 8028, 8151, 8157) with the Subaru 8.2-m telescope (+ FOCAS) on June 27.6 UT. SN 2002ic is still bright (R ∼ 19) and still shows a strong Hα emission line (FWHM ∼ 6 nm) and a weaker Hβ emission line (Hα/Hβ flux ratio ∼ 16). A broad and strong Ca II triplet emission is also seen. The spectrum has an apparent similarity to that of the peculiar type-IIn supernova 1997cy between ∼ 100 and 250 days after the explosion (Turatto et al. 2000, Ap.J. 534, L57), and this is consistent with the metamorphosis observed in the earlier phase (IAUC 8151). The depolarization across the Hα emission is distinct, which indicates that SN 2002ic has a considerable intrinsic polarization (p > 1 percent) at some wavelength regions. The spectrum of SN 2002ic is dominated by the interaction region and appears similar to a type-IIn supernova spectrum.”

2003 July 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green